HUBBLE SPACE TELESCOPE DAILY REPORT #2223 PERIOD COVERED: 0000Z (UTC) 10/05/98 - 0000Z (UTC) 10/06/98 Daily Status Report as of 279/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled. 1.2 Completed STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8075 (HDF-S Campaign) The WF/PC-2, The Space Telescope Imaging Spectrograph (MA2 and CCD) and the Near Infrared Camera (Cameras One, Two and Three) all were used to make deep observations of the Hubble Deep Fields- South (HDFS-22333). This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: The proposals detailed in today's report support the beginning of the second and final week of the Hubble Deep Field - South (HDF-S) campaign. All HDF-S proposals completed nominally today. Using ROP SR-1, the SSR EDAC error counter was reset three times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH